Bond C-C C-O (in CO2) O-O (in O2) O-H C-H Avg Bond Strength (in kJ/mol) 350 804 498 464 415The table above gives the average bond strength for a variety of chemical bonds. Use these values to calculate:

-The energy change (in kJ) for the combustion of 105.6 grams of propane?

-The energy change (in kJ) for the combustion of 2.65 grams of ethane using 13.70 grams of oxygen?
Answer #1

Ans 1

Moles of propane = mass/molecular weight

= 105.6g / 44.1g/mol

= 2.395 mol

Combustion of propane

C3H8 + 5O2 = 3CO2 + 4 H2O

Enthalpy change

H = sum of Enthalpy change in bond formation - sum of Enthalpy change in bond broken

H = 2*(C--C) + 8*(C--H) + 5*(O=O) - 6*(C=O) - 8*(O--H)

= 2*(350) + 8*(415) + 5*(498) - 6*(804) - 8*(464)

= - 2026 kJ/mol

energy change (in kJ) for the combustion of 105.6 grams of propane

= - 2026 kJ/mol x 2.395 mol

= - 4852.27 kJ

Ans 2

Moles of C2H6 = 2.65g / 30.07g/mol

= 0.0881 mol

Moles of O2 = 13.70g / 32g/mol

= 0.4281 mol

2C2H6 + 7O2 = 4CO2 + 6H2O

From the stoichiometry of the reaction

2 mol C2H6 required = 7 mol O2

0.0881 mol C2H6 required = 7*0.0881/2 = 0.308 mol O2

C2H6 limiting reactant

H = 2*(C--C) + 12*(C--H) + 7*(O=O) - 8*(C=O) - 12*(O--H)

= 2*(350) + 12*(415) + 7*(498) - 8*(804) - 12*(464)

= - 2834 kJ/mol x 0.0881 mol C2H6

= - 249.75 kJ